OpenStudy (anonymous):

an airplane flies 240 miles against the wind in the same amount of time it takes to fly 360 miles with the wind. the speed of plane is still air is 30 mph more than 4 times the speed of the wind. Find the speed of the plane inn still air

OpenStudy (perl):

let x = speed of plane in still air y= wind speed x = 4*y + 30 the distance equations give us 240 = ( x - y ) *t 360 = (x +y)*t t is the same in both equations 240 / ( x - y) = 360 / ( x + y ) now plug in x = 4y + 30 for x